OpenStudy (anonymous):

HELP I DON'T HAVE A CALCULATOR GOOD ENOUGH FOR THIS!! Reba is using the TVM Solver on her graphing calculator as shown below to determine the future value of $2400 after it has earned compound interest for a certain number of years. According to what Reba has entered into the TVM Solver, for how many years will the $2400 earn compound interest? A. 12 B. 3 C. 6 D. 4

OpenStudy (anonymous):

What is her initial value?

OpenStudy (anonymous):

I think we are missing a couple of important pieces of info here

OpenStudy (anonymous):

sorry for the wait there is a picture saying : N=4 I%=6 PV=-2400 PMT=0 FV= P/Y=1 C/Y=12

OpenStudy (anonymous):

which I don't quite understand that

OpenStudy (anonymous):

N usually stands for the amount of years

OpenStudy (anonymous):

thank you very much!!

OpenStudy (anonymous):

Its slightly confusing though since PV =-2400 and FV=2400 That basically means Reba earned nothing ...but either way N in this case seems to be referring to the number of years

OpenStudy (anonymous):

And thank you for taking the time to explain:)! I I understand it much more now.